OpenStudy (anonymous):

abcd is a rectangle. find the length of each diagonal. ac = 2(x-3) bd = x + 5

OpenStudy (anonymous):

The diagonals (ac and bd) of a rectangle are of equal length so....

OpenStudy (anonymous):

SOOOO THEN WHATS THE ASNWER ???

OpenStudy (anonymous):

No need to shout, I thought you might be able to manage it yourself:-) If 2x+3 = x+5 -> x = 2.

OpenStudy (anonymous):

wasnt shoutinq , sorry had accidentl y pressed the caps button

OpenStudy (anonymous):

No prob.

OpenStudy (anonymous):

but that isnt the correct answer

OpenStudy (anonymous):

ooOO nooo im sooo sorry u answered my last question , the question i needed answered was abcd is a rectangle. find the length of each diagonal.ac=2(5a+1) bd=2(a+1). im sorry lol

OpenStudy (anonymous):

oops, apologies, dunno where I got 2x+3 from, so 2(x-3) = x+5 -> x = 11

OpenStudy (anonymous):

Your equations produce 8a = 0 so your diagonals are 2.

OpenStudy (anonymous):

6 Made_ForThis: find the value of x. rt=4x+10 sw=5x-20